OpenStudy (lovelycharm):

Select the best possible first step to solving the system by first eliminating the x variable. 3x − 9y = 6 2x − 11y = 6

OpenStudy (solomonzelman):

multiply the first equation times 2, and multiply the second equation times -3. Then, add the equations together.

OpenStudy (anonymous):

^^^^^^

OpenStudy (lovelycharm):

6x-18y=12 -6x+33y=-18

OpenStudy (solomonzelman):

yes, then add the equations:

OpenStudy (lovelycharm):

15=30

OpenStudy (solomonzelman):